OpenStudy (anonymous):

2(8u+2)=3(2u-7) 16u+4=6u-21 16u+4-6u=6u-21-6u 10u+4=(-21) 10u+4-4=(-21)-4 10u=-25 10u/10=-25/10 u=-2.5 how do i justify it

OpenStudy (anonymous):

u just did

jimthompson5910 (jim_thompson5910):

2(8u+2)=3(2u-7) 2(8(-2.5)+2)=3(2(-2.5)-7) ... replace ALL copies of 'u' with -2.5 (since u = -2.5) 2(-20+2)=3(-5-7) 2(-18)=3(-12) -36=-36 Since the last equation is true, this verifies the answer

OpenStudy (anonymous):

it doesn't get any more justified than that

jimthompson5910 (jim_thompson5910):

Unfortunately @sonofa_nh it doesn't always justify since you could be introducing extraneous solutions that do not work

OpenStudy (anonymous):

well the paper says to put it in seven lines a. b. c. d. e. f. g.

jimthompson5910 (jim_thompson5910):

does the idea make sense coraline?

OpenStudy (anonymous):

eh...........no

jimthompson5910 (jim_thompson5910):

I'm replacing all copies of the variable 'u' with -2.5 (since u is -2.5) and then evaluating like normal

jimthompson5910 (jim_thompson5910):

at the end, I get a true equation, so u = -2.5 works

OpenStudy (anonymous):

but how do i put it into seven lines?

jimthompson5910 (jim_thompson5910):

Simple example: solve x+2=4 solution is x = 2 check: x+2 = 4 2+2 = 4 4 = 4 ... works

jimthompson5910 (jim_thompson5910):

break up one line into two to make more lines

jimthompson5910 (jim_thompson5910):

but what I wrote should be enough

OpenStudy (anonymous):

ok i think i get it so i break up each line???

jimthompson5910 (jim_thompson5910):

just enough to get 7 lines

jimthompson5910 (jim_thompson5910):

a. 2(8u+2)=3(2u-7) b. 2(8(-2.5)+2)=3(2(-2.5)-7) c. 2(-20+2)=3(2(-2.5)-7) d. 2(-20+2)=3(-5-7) f. 2(-18)=3(-5-7) e. 2(-18)=3(-12) g. -36=-36

jimthompson5910 (jim_thompson5910):

I'm basically streeeeetching it out to make it fit

OpenStudy (anonymous):

thanks